OpenStudy (logan13):

The East-Side soccer team has a cylindrical water cooler that is half full. The diameter of the cooler is 12 in. and the height is 15 in. The team’s drinking cups hold 20.14 in3. How many drinking cups can be filled from the cooler? Use 3.14 to approximate pi, and round your answer to the nearest whole number. A. 14 B. 42 C. 84 D. 169

563blackghost (563blackghost):

You would first find the volume of the entire cooler and then divide by 2... \(\huge{V=\frac{ \pi r^{2}h}{2}}\) So we input what we know... \(\huge\color{purple}{V=\frac{ 3.14(6^{2})15}{2}}\) We have \(\large{6}\) as the radius because that is half of \(\large{12}\)... So we simplify to find the volume of the liquid....

563blackghost (563blackghost):

Were you able to get the Volume?